### v1

# Towards Imperceptible JPEG Deep Hiding: Multi-range Representations-driven Adversarial Stego Generation

Abstract
Deep hiding has been exploring the hiding capability of deep learning-based models, aiming to conceal image-level messages into cover images and reveal them from generated stego images. Existing schemes are easily detected by steganalyzers due to their large payloads and their limitation to feature extraction based solely on either pure convolution or pure transformer operators within a single range, as well as pixel-level loss constraints. To address the issue, in this paper, we introduce generation-based adversarial attacks into color JPEG image deep hiding and propose a multi-range representations-driven adversarial stego generation framework called MRAG from a steganalysis perspective. Specifically, we integrate the local-range neighbor reception characteristic of the convolution and the global-range dependency modeling of the transformer to construct MRAG. Meanwhile, we use the transformed images obtained through coarse-grained and fine-grained frequency decomposition as inputs, introducing multi-grained information. Furthermore, a features angle-norm disentanglement loss is designed to constrain the generated stegos closer to covers in the angle and norm space of the steganalyzer’s classified features. Consequently, small yet effective adversarial perturbations can be injected into the process of generating stegos, ensuring that stegos maintain favorable secret restorability and imperceptibility. Extensive experiments demonstrate that MRAG can achieve state-of-the-art performance.
1 Introduction
Image steganography is the art and science of sheltering confidential information in publicly available cover images to obtain stego ones, from which secret information can be revealed, and resulting stegos should be as unsuspicious as possible Provos and Honeyman (2003). It usually has three primary metrics Pevný and Fridrich (2008): the capacity refers to embedded payloads, the visual imperceptibility regarding the cover and stego being indistinguishable by naked eyes, and the steganalysis imperceptibility against the steganalyzers. Traditional steganographic methods and earlier deep learning-based methods maintain the bit-level payloads, typically setting it below 0.4 bits per pixel (BPP), with traditional methods going up to 1 BPP at most, and deep learning-based methods reaching up to 4.5 BPP Lu et al. (2021a). In light of this, Baluja Baluja (2017) delves deeper into the potential of steganographic models based on deep learning, successfully hiding a spatial RGB image into another one of the same size, known as deep hiding. Three natural advantages of deep hiding are also pointed out: 1) Since the type of hidden secret information is an image with high inherent correlations, the difficulty of network learning can be appropriately reduced. 2) The secret itself has a certain degree of fault tolerance. 3) The magnitude of the payload can be up to 24 BPP with a 1:1 ratio between the hidden secret image and the cover image. The success of the pioneering work, along with three natural advantages, paves the way for the flourishing of deep hiding.
The act of hiding inevitably yields embedded artifacts within the stegos. Whether artifacts are discovered usually depends on three factors Baluja (2020): the hidden capacity, the imperceptibility of the algorithm, and the texture of the image itself. The likelihood of the hiding behavior being discovered is negatively correlated with the latter two factors, and positively correlated with the capacity. In this sense, the capacity advantage of deep hiding also makes it more susceptible to detection, turning it into a disadvantage. Deep hiding with spatial RGB images as covers, abbreviated as DH-Spatial, has made significant progress recently Zhang et al. (2020); Lu et al. (2021a); Xu et al. (2022). However, the latest work, EFDR Yang and Liao (2023), points out that DH-Spatial methods perform poorly when the cover is the ubiquitous lossy image format - color JPEG image, resulting in visible color distortion, blocking, blurring, and banding artifacts in the generated JPEG stego. Similar visible artifacts also appear in the restored secret revealed from the JPEG stego, as shown in Fig.1 (a). Although EFDR achieves acceptable visual imperceptibility and restorability, its steganalysis imperceptibility remains unsatisfactory, as depicted in Fig.1 (b). The common shortcomings of these schemes are that they rely solely on single-range feature extraction using either pure convolution or pure transformer operators, and only consider pixel-level loss constraints.
A steganalyzer is a binary classifier that distinguishes between cover and stego images. With the development of deep learning, the way of features extraction has changed from manual to automatic extraction by networks and has become mainstream Wei et al. (2024); Hu and Wang (2024). Deep learning-based classifiers are found to be vulnerable to adversarial examples that can cause misclassification, which brings the prospect of resisting steganalyzers based on deep learning Zhang et al. (2018b). Reference Szegedy et al. (2014) puts forward that the adversarial example can be viewed as “accidental steganography”. They are achieved by adding low amplitude signals to an image for adversarial attacks or secret embedding and extraction. Analogously, deep hiding and generation-based adversarial attacks employ an image-to-image generative architecture to learn the image transformation task from the input image to the generated image Liu and Hsieh (2019). Therefore, in this paper, we attempt to introduce generation-based adversarial attacks into color JPEG image deep hiding, exploring the potential of generated color JPEG stego images in terms of secret restorability and imperceptibility, as illustrated in Fig.1 (c).
For the steganalyzer, to better capture subtle changes, it is essential to consider both local-range and global-range information Wei et al. (2024). This in turn inspires us to integrate the local-range neighbor reception characteristic of convolution and the global-range dependency modeling of transformer, to construct a multi-range representations-driven adversarial stego generation framework called MRAG from a steganalysis perspective. Specifically, MRAG comprises four key components: the invertible local-range branch, the invertible global-range branch, the adaptive fusion module, and the surrogate steganalyzer. We also design the coarse-grained frequency decomposition (CFD) and the fine-grained frequency decomposition (FFD) to transform the input. According to the dimension characteristic of the obtained transformed images, the transformed images are used as the input for the local and global branches, respectively, naturally introducing multi-grained frequency information. Moreover, a features angle-norm disentanglement loss is designed to disentangle the surrogate steganalyzer’s classified features into the angle space associated with inter-class semantic differences and the norm space related to intra-class confidence, respectively. It can constrain the generated stegos closer to covers by pulling together the angle between the classified features of cover and stego, while keeping the norm changeless. With the help of features angle-norm disentanglement loss, small yet effective adversarial perturbations can be injected based on the multi-range representations extracted by local and global branches, ensuring that stegos maintain favorable secret restorability and imperceptibility. Our contributions can be summarized as follows:
We propose a multi-range representations-driven adversarial stego generation framework, MRAG, to generate color JPEG stego images with favorable secret restorability and imperceptibility. To our knowledge, this is the first to focus on the steganalysis imperceptibility issue of color JPEG image deep hiding.
We integrate the local-range neighbor reception of the convolution and the global-range dependency modeling of the transformer to extract multi-range representations from a steganalysis perspective, injecting adversarial perturbations based on these representations.
A features angle-norm disentanglement loss is designed to constrain the angle and norm distances between the steganalyzer’s classified features of the cover and the stego, in which the stego can be closer to the cover.
Experiments validate the effectiveness of MRAG, demonstrating that MRAG establishes a new state-of-the-art benchmark in color JPEG image deep hiding.
2 Related Work
In this section, we summarize advancements in deep hiding and briefly introduce generation-based adversarial attacks.
2.1 Deep hiding
Seminal work Baluja (2017) designs a deep hiding model that contains a secret pre-processing module, a hiding module, and a revealing module. UDH Zhang et al. (2020) is a cover-agnostic deep hiding model, which discards the pre-processing structure and only encodes secret images in the hiding sub-network. ISN Lu et al. (2021a) treats embedding and extraction as an inverse problem and utilizes the reversibility of invertible neural networks to address this question through its forward and backward processes. HiNet Jing et al. (2021) combines invertible neural networks with the wavelet domain and introduces low-frequency wavelet loss. RIIS Xu et al. (2022), a robust deep hiding model, takes various distortions into consideration, which shows acceptable robustness. The cover types and operators involved in the aforementioned works are always spatial RGB images and pure convolution. EFDR Yang and Liao (2023) is the first to research hiding image-level information in color JPEG images, focusing on mining fine-grained DCT representations using the pure transformer operator. Although advances in deep hiding have been remarkable, existing schemes are limited to single-range feature extraction, relying solely on either pure convolution or pure transformer operators. Moreover, these schemes are constrained only by pixel-level loss, leading to poor imperceptibility. It should be noted that, in this paper, we also perform deep hiding in color JPEG images, the ubiquitous lossy format images in daily life, and pay attention to its steganalysis imperceptibility issue first.
2.2 Generation-based adversarial attacks
Reference Szegedy et al. (2014) unveils that deep learning-based classifiers are susceptible to adversarial examples. Adding tiny perturbations to the input can cause misclassification with high confidence. Adversarial attacks can mainly be divided into two categories: gradient-based attacks and generation-based attacks Macas et al. (2024). Generation-based attacks usually utilize an image-to-image generative architecture to learn the mapping from input images to adversarial ones. Once the adversarial generator is trained well, only one forward propagation is needed to produce plenty of adversarial examples quickly. For example, AdvGAN Xiao et al. (2018), a representative of generation-based attacks, consists of an adversarial generator, a discriminator that differentiates between the original and the adversarial example, and a surrogate classifier intended to be attacked. However, SSAE Lu et al. (2021b) demonstrates that the discriminator is redundant if perturbations are tiny enough. Therefore, our MRAG also discards the discriminator to simplify the model.
3 Proposed Framework
3.1 Overview
The pipeline of MRAG consists of four key components: the invertible local-range branch, the invertible global-range branch, the adaptive fusion module, and the surrogate steganalyzer, as shown in Fig. 2. Since we construct MRAG utilizing invertible neural networks as much as possible to ensure reversibility, it can implement forward-hiding and backward-revealing processes with shared parameters. The other image conversion operations and network structures are also reversible and symmetric. Conversion operations contain the coarse-grained frequency decomposition (CFD), the fine-grained frequency decomposition (FFD), and the inverse FFD (IFFD). The input of MRAG’s forward-hiding process is the color JPEG cover image and the spatial RGB secret image . The input is converted into coarse-grained and fine-grained frequency maps through CFD and CFDFFD, respectively. According to the dimension characteristic of the obtained frequency maps, they are fed into the local and global branches’ forward processes, respectively. Then, the global branch’s output passes through the IFFD and is concatenated with the local branch’s output. Afterward, their concatenation is fed into the adaptive fusion module’s forward process to obtain the JPEG stego image and the redundant information , an all-zero matrix. MRAG’s backward-revealing process takes the color JPEG stego image and as input to reveal the secret image. For effectively inducing the surrogate steganalyzer to misclassify the cover and stego , we conduct the features angle-norm disentanglement and pull together the angles between the classified features of cover and the classified features of stego while keeping their norm values changeless.
3.2 MRAG Model
Backbone structure.
Given that invertible neural networks can preserve as much detail as possible due to the theoretically lossless information characteristic Dinh et al. (2014, 2017); Gilbert et al. (2017), we utilize invertible neural networks to construct our MRAG to the greatest extent. As shown in Fig. 2, several key components of MRAG, including the invertible local-range branch, the invertible global-range branch, and the adaptive fusion module, all employ invertible neural networks. For the basic unit of invertible neural networks, the affine coupling layer, the input needs to be divided into two parts, and , along the channel axis. Following reference Xiao et al. (2020), we use an additive affine transformation for the upper part, , and employ an enhanced affine transformation for the lower part, . To be specific, the forward process of the affine coupling layer can be formulated as
In Eq. (1), , , and can be arbitrary network structures. and are the Hadamard product and Exponential function, respectively. Note that is omitted in Fig. 2. Correspondingly, the inverse process is
In the following, we will provide a detailed description of the four key components of MRAG.
Invertible local-range branch.
Convolution performs excellently in image-related tasks and has nice local neighbor reception characteristics Krizhevsky et al. (2012). Here, , , and in the invertible local-range branch is implemented by a 5-layer dense block with a convolution kernel Huang et al. (2017). In the dense block, the output feature maps of each layer are passed to all subsequent layers, greatly enhancing the ability for representation sharing and learning. For the invertible local-range branch, the input is coarse-grained frequency maps through CFD. CFD includes the Huffman decoding of the color JPEG cover image and the DCT transformation of the spatial RGB secret image.
Invertible global-range branch.
To further enhance representation and global dependency modeling ability, in the invertible global-range branch, , , and adopt a standard 1-layer Pre-LN transformer Dosovitskiy et al. (2021). This Pre-LN transformer applies the layer normalization before the multi-head self-attention and the multi-layer perceptron for a more stable gradient distribution. For the invertible global-range branch, the input is the fine-grained frequency maps obtained through CFDFFD. FFD includes the sub-band gathering and the zigzag scanning of the coarse-grained frequency maps. The sub-band gathering procedure rearranges coarse-grained frequency maps to place the frequency components of the same sub-band on the same channel, making frequency components more compact. The zigzag scanning procedure further rearranges them along the channel axis to group similar frequency sub-bands together, exploiting the structural information in the JPEG transformation.
Adaptive fusion module.
Our designed adaptive fusion module consists of a pair of convolutions with channel attention and a series of affine coupling layers with an invertible convolution. , , and in the adaptive fusion module also employ a 5-layer dense block. The convolution with channel attention designed in reference Tan et al. (2022) is used to adaptively extract the most informative frequency representations. The invertible convolution is a kind of convolution commonly used for invertible neural networks. Adding an invertible convolution before each affine coupling layer can increase the dependencies between representations, allowing for more effective information fusion. For the adaptive fusion module, the input is the concatenation of representations extracted from the invertible local-range branch and representations extracted from the invertible global-range branch after undergoing IFFD. IFFD is used for dimension alignment.
Surrogate steganalyzer.
In our MRAG, the surrogate steganalyzer can be any well-trained JPEG steganalyzer. Here, we select three open-source steganalyzers: UCNet Wei et al. (2022), EWNet Su et al. (2021), and DBS2Net Hu and Wang (2024). UCNet is the sole steganalyzer tailored for color JPEG images. The others are straightforwardly based on existing steganalyzers for gray-scale JPEG images or on some well-known models in computer vision. EWNet and DBS2Net are gray-scale JPEG steganalyzers. Here, following the setting in reference Wei et al. (2022), we adjust the number of input channels in the first convolution layer to 3, thereby enabling them to detect color images.
Method | COCO | ImageNet | BOSSBase | |||||||||
PSNR | SSIM | APD | LPIPS | PSNR | SSIM | APD | LPIPS | PSNR | SSIM | APD | LPIPS | |
ISN+ Lu et al. (2021a) | 19.85 | .662 | 17.52 | .344 | 19.81 | .643 | 18.98 | .361 | 22.55 | .731 | 14.60 | .345 |
RIIS Xu et al. (2022) | 24.98 | .798 | 9.28 | .169 | 24.97 | .781 | 9.80 | .178 | 29.20 | .864 | 5.84 | .143 |
EFDR Yang and Liao (2023) | 33.95 | .918 | 4.31 | .028 | 34.05 | .919 | 4.37 | .027 | 36.43 | .921 | 3.22 | .032 |
MRAG (UCNet) | 40.11 | .983 | 1.88 | .004 | 39.85 | .980 | 2.04 | .004 | 41.54 | .975 | 1.66 | .005 |
MRAG (EWNet) | 41.67 | .988 | 1.62 | .001 | 41.20 | .985 | 1.80 | .001 | 42.32 | .978 | 1.55 | .002 |
MRAG (DBS2Net) | 37.14 | .978 | 2.78 | .006 | 37.00 | .973 | 2.90 | .005 | 39.49 | .969 | 2.15 | .007 |
Method | COCO | ImageNet | BOSSBase | |||||||||
PSNR | SSIM | APD | LPIPS | PSNR | SSIM | APD | LPIPS | PSNR | SSIM | APD | LPIPS | |
ISN+ Lu et al. (2021a) | 24.09 | .828 | 9.96 | .208 | 24.08 | .823 | 10.66 | .213 | 27.12 | .830 | 7.80 | .239 |
RIIS Xu et al. (2022) | 28.25 | .897 | 5.63 | .098 | 28.56 | .895 | 5.86 | .104 | 32.76 | .921 | 3.93 | .102 |
EFDR Yang and Liao (2023) | 32.22 | .927 | 4.71 | .025 | 32.04 | .924 | 4.93 | .027 | 36.01 | .944 | 3.12 | .027 |
MRAG (UCNet) | 31.68 | .920 | 5.01 | .028 | 31.53 | .916 | 5.24 | .031 | 35.60 | .940 | 3.27 | .031 |
MRAG (EWNet) | 32.78 | .936 | 4.39 | .023 | 32.62 | .934 | 4.58 | .025 | 36.37 | .950 | 2.98 | .026 |
MRAG (DBS2Net) | 32.35 | .930 | 4.65 | .022 | 32.16 | .927 | 4.86 | .024 | 35.92 | .945 | 3.16 | .026 |
3.3 Loss Function
Features angle-norm disentanglement loss.
Several works Liu et al. (2018); Chen et al. (2020) conjecture that classified features learned by classifiers can naturally disentangle into angle and norm distances related to inter-class semantic differences and intra-class confidence, respectively. The steganalyzer can be viewed as a binary classifier. The classified features learned by the steganalyzer are crucial for distinguishing between covers and stegos. Inspired by these facts, we introduce this loss to constrain the generated stegos closer to covers in the angle and norm space of the steganalyzer’s classified features, thereby enhancing the steganalysis imperceptibility. For fooling the surrogate steganalyzer, we focus on pulling together the angles between the classified features of cover and the classified features of stego . Additionally, to minimize the injected adversarial perturbations, we keep the norm values of and changeless. Thus, we can define the features angle-norm disentanglement loss as follows
where indicates the cosine similarity. measures the similarity between two feature vectors by the cosine value of the angle between them. A larger value indicates a smaller angle, and hence a higher similarity. is the mean squared error. and denote the angle loss and norm loss of the classified features, respectively.
Total loss.
The total loss consists of three parts: the hiding loss to ensure visual imperceptibility, the revealing loss to ensure secret restorability, and the features angle-norm disentanglement loss to guarantee steganalysis imperceptibility. is equal to the sum of these three losses, expressed as follows
In Eq. (4), represents the actual restored secret image.
Method | COCO | ImageNet | BOSSBase | ||||||
ACC_U | ACC_E | ACC_D | ACC_U | ACC_E | ACC_D | ACC_U | ACC_E | ACC_D | |
EFDR Yang and Liao (2023) | 92.73 | 75.68 | 99.90 | 92.85 | 75.65 | 99.90 | 87.65 | 73.70 | 99.98 |
MRAG (UCNet) | 50.05 | 51.28 | 74.53 | 50.00 | 51.68 | 78.83 | 50.00 | 50.55 | 63.28 |
MRAG (EWNet) | 58.53 | 60.48 | 71.60 | 60.73 | 59.25 | 74.65 | 51.20 | 55.00 | 59.05 |
MRAG (DBS2Net) | 53.35 | 51.88 | 50.13 | 53.50 | 51.18 | 50.30 | 51.25 | 50.43 | 50.13 |
(a) UCNet_EFDR
(c) EWNet_EFDR
(e) DBS2Net_EFDR
(b) UCNet_MRAG
(d) EWNet_MRAG
(f) DBS2Net_MRAG
The Number of Layer | Secret Restorability | Visual Imperceptibility | Steganalysis Imperceptibility | #Param () | ||||
PSNR | SSIM | PSNR | SSIM | ACC_U | ACC_E | ACC_D | ||
1 | 41.59 | .987 | 33.21 | .941 | 50.05 | 52.48 | 75.88 | 14.54 |
2 | 40.56 | .984 | 32.60 | .934 | 50.08 | 52.85 | 98.90 | 14.73 |
3 | 40.11 | .983 | 31.68 | .920 | 50.05 | 51.28 | 74.53 | 14.91 |
4 | 39.72 | .983 | 31.80 | .923 | 50.05 | 64.23 | 80.30 | 15.10 |
Local Range | Global Range | Secret Restorability | Visual Imperceptibility | Steganalysis Imperceptibility | ||||
PSNR | SSIM | PSNR | SSIM | ACC_U | ACC_E | ACC_D | ||
✗ | ✔ | 26.23 | .786 | 31.95 | .924 | 50.03 | 51.15 | 67.75 |
✔ | ✗ | 36.20 | .965 | 31.27 | .908 | 50.18 | 51.10 | 85.13 |
✔ | ✔ | 40.11 | .983 | 31.68 | .920 | 50.05 | 51.28 | 74.53 |
4 Experiment
4.1 Experimental Settings
Datasets.
We employ three image datasets in our implementation, including COCO Lin et al. (2014), ImageNet Russakovsky et al. (2015), which are widely used in various deep learning tasks, and a standard steganographic dataset BOSSBase Bas et al. (2011). We utilize 5000, 1000, and 2000 cover-secret pairs from COCO for training, validation, and testing, respectively. 2000 cover-secret pairs from ImageNet and BOSSBase, respectively, are used only for testing. Additional 10000 cover-stego pairs, randomly mixed from three datasets, are used for training the steganalyzer. The stego images are generated by HRJS Yang et al. (2023), a bit-level JPEG steganographic method based on deep learning. All images are center-cropped to . The compression quality factor adopted by the cover is 75, the most commonly used in JPEG steganography.
Baselines.
We take several advanced and representative methods as our baselines, including ISN Lu et al. (2021a), a DH-Spatial method, RIIS Xu et al. (2022), a robust DH-Spatial method, and EFDR Yang and Liao (2023), which is currently the only customized deep hiding method with color JPEG images as covers. Note that during the training phase of ISN, we add an additional JPEG compression distortion layer, referred to as ISN+ for distinction.
Evaluation metrics.
To measure the secret restorability and visual imperceptibility, we adopt peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) Wang et al. (2004), average pixel discrepancy (APD), and learned perceptual image patch similarity (LPIPS) Zhang et al. (2018a). PSNR/SSIM with larger values and APD/LPIPS with smaller values indicate higher image quality. For steganalysis imperceptibility, we utilize the detection accuracy rate (ACC) as the metric. The ACC is closer to 50%, i.e., random guess, indicating superior performance. ACC_U, ACC_E, and ACC_D denote the detection accuracy rate of UCNet, EWNet, and DBS2Net, respectively.
Implementation details.
Before training MRAG, the surrogate steganalyzer is pre-trained 50 epochs on 5000 mixed cover-stego pairs with the parameters following the default settings. Subsequently, the pre-trained surrogate steganalyzer’s weights are frozen, and we train MRAG for 200 epochs on 5000 cover-secret pairs from COCO. We adopt the Adam optimizer with , , and , and set the learning rate to . The mini-batch size is 4, containing two randomly selected cover-secret pairs. Note that the steganalyzer used for the steganalysis imperceptibility evaluation is trained on another 5000 mixed cover-stego pairs. All experiments are implemented with PyTorch and on the NVIDIA GeForce RTX 4090 GPU.
4.2 Comparison Results
In this subsection, to verify the performance of MRAG in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility, we conduct extensive experiments on three datasets and three open-source JPEG steganalyzers.
Secret restorability.
Table 1 reports the numerical comparisons of our MRAG with existing baselines. It can be observed that MRAG significantly outperforms these baselines in terms of secret restorability. Compared with EFDR, MRAG can supply at least 2.95 dB PSNR advancement. Similar performance improvements can also be observed in SSIM, APD, and LPIPS. Visualization comparisons are shown in Fig. 3. We can observe that the secret images revealed by MRAG are visually pleasing with minimal visual artifacts. In contrast, the DH-Spatial methods ISN+ and RIIS yield visible color distortion, blocking, blurring, and banding artifacts in the revealed secret images. It is noteworthy that although our model only uses the COCO for training, it still provides excellent results on the other two datasets, which indicates that our model has a favorable generalization ability.
Visual imperceptibility.
Table 2 and Fig. 4 report the numerical comparisons and visualization comparisons of our MRAG with ISN+, RIIS, and EFDR in terms of visual imperceptibility. We can observe that MRAG and EFDR have comparable performance, while the stego images generated by ISN+ and RIIS still contain many visible artifacts, making it difficult for them to achieve visual imperceptibility.
Steganalysis imperceptibility.
Steganalysis imperceptibility, an essential metric, evaluates the statistical imperceptible to resist steganalyzers. It achieves classification by distinguishing the changes in statistical features caused by hiding. Table 3 shows the numerical comparisons of our MRAG with EFDR. It can be observed that MRAG significantly outperforms EFDR in terms of steganalysis imperceptibility. Compared to EFDR, MRAG can provide at least 15.2% improvement, with most classification results approaching 50%. Moreover, we visualize the distribution of the classified features learned by three steganalyzers using t-distributed stochastic neighbor embedding (t-SNE) in Fig. 5. We can find that the distribution of covers’ classified features (green points) and stegos’ classified features (red points) obtained by EFDR can be well distinguished, while the distribution of the classified features obtained by the proposed MRAG is intermingled and difficult to distinguish. This also confirms the superior steganalysis imperceptibility of MRAG.
4.3 Ablation Study
Here, we mainly discuss the impact of the number of the affine coupling layer and multi-range branches.
Impact of the number of the affine coupling layer.
For the convenience of discussion, we fix the number of the affine coupling layer in the invertible local-range branch and the invertible global-range branch to 1, and focus on investigating the impact of the number of the affine coupling layer in the adaptive fusion module on performance. As reported in Table 4, we can observe that the steganalysis imperceptibility performance, the most concerned metric, is optimal when the number of layers is 3. Thus, in our implementation, the number of the affine coupling layer in the adaptive fusion module defaults to 3 unless otherwise specified.
Effectiveness of multi-range branches.
Also, we verify the performance effects of the invertible local-range branch and the invertible global-range branch, as shown in Table 5. It can be observed that compared to any single-range branch, our designed multi-range branches can significantly boost the performance of secret restorability and steganalysis imperceptibility while keeping visual imperceptibility.
5 Conclusion
In this work, we attempt to introduce generation-based adversarial attacks into color JPEG image deep hiding, exploring the potential of generated color JPEG stego images in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility. To the best of our knowledge, this is the first attempt to focus on the steganalysis imperceptibility issue of color JPEG image deep hiding. A multi-range representations-driven adversarial stego generation framework, MRAG, is proposed from a steganalysis perspective. Besides, a features angle-norm disentanglement loss is designed to constrain the generated stegos closer to covers in the angle and norm space of the steganalyzer’s classified features. Comprehensive experiments verify the effectiveness of MRAG and its superior performance. We believe the success of MRAG can further promote the development of color JPEG image deep hiding and provide some inspiration for further research on steganalysis imperceptibility.